HIGH-RESOLUTION IR ACTION SPECTRUM OF C_{2}H_{2}^{+}

\underline{SABRINA GÄRTNER}, JÜRGEN KRIEG, OSKAR ASVANY and STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln.

The method of Laser-Induced-Reactions (LIR), which is described in detail in Schlemmer et al. 2002^{a}, is used to obtain high-resolution infrared spectra of molecular ions. Here, the endothermic reaction C_{2}H_{2}^{+} + H_{2} → C_{2}H_{3}^{+} + H is promoted by ro-vibrational excitation of the parent ion. A spectrum of the ν_{3} stretching vibration of C_{2}H_{2}^{+} is recorded by variation of the wavelength of a home-build OPO (Optical Parametric Oscillator) operating in the 3 micron region. The experiments are carried out in a low temperature 22-pole ion trap where several hundred cold, mass selected parent ions are stored. Typical linewidths of the action spectra are 3 × 10^{−3} cm^{−1}. First spectra for C_{2}H_{2}^{+} and their analysis will be presented. Other possible applications of LIR spectroscopy will be discussed.

aS. Schlemmer, E. Lescop, J. von Richthofen, D. Gerlich, and M. Smith, Laser Induced Reactions in a 22-Pole Ion Trap:

C_{2}H_{2}^{+} + H_{2} → C_{2}H_{3}^{+} + H, J. Chem. Phys. \underline{117}(2068-2075), 2002